OpenStudy (anonymous):

When [equation] is divided by (x-1) the remainder is 7. When it is divided by (x+1) the remainder is 3. Determine the values of a &b . Can somebody please just explain the method... thanks:)

OpenStudy (anonymous):

\[\huge x^4-4x^3+ax^2+bx+1 \]

OpenStudy (anonymous):

Have you tried synthetic division?

OpenStudy (anonymous):

but there are 3 variables .. :S

OpenStudy (anonymous):

Yeah but you know a lot of algebra to deal with them.

OpenStudy (anonymous):

Hmmm, I'm not exactly sure what the trick for this is.

OpenStudy (anonymous):

\[ \large x^4-4x^3+ax^2+bx+1=(x-1)(cx^3+dx^2+ex+f)+7 \]

OpenStudy (anonymous):

@ burhan101 use factor and remainder theorem

OpenStudy (anonymous):

when f(x) is divided by (x-a) the remainder is given by f(a)

OpenStudy (anonymous):

thus here f(x) =x^4−4x^3+ax^2+bx+1 f(1)=7 and f(-1)=3 u will get two equation in two unknowns and u can solev them to know a and b

OpenStudy (anonymous):

thus a+b=9 and a-b -3 thus solving above two eq we have a=3 and b=6